Grade 12th passIIT JEE Entrance Exam

given that p=q=r. if pq=r then the angle between p and r is theta1 if pq=0 then the angle b/w p and r is theta2 . what is the relation b/w theta1 and theta2 1)theta1=2 2) theta1 =theta2/2 3) theta1 =2theta2 4)none

Deeksha
9 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
 Magnitude Sum of Two vectors Vector P and Vector q is given by
r^2 =p^2+q^2+2pqcos($)   __________________________(i)
here $ indicates angle between them.
We are given 
r=p+q
so Squaring  both the sides we  get
r^2=p^2+q^2+2pq    ______________________(ii)
From (i) and (ii) we can write
p^2+q^2+2pqcos($)=p^2+q^2+2pq
therefore 2pq=2pqcos($)
cos($)=1
therefore $=0
So angle between p vector and q vector is 0 degree.